OpenStudy (anonymous):

Let L be the tangent to f(x)=x^.5 at any point. Show that the x-intercept of the tangent to the curve at the point (h,h^.5) is -h. Please help!!!

OpenStudy (anonymous):

h^.5=1/(2 sqrt(h)) h + b b = sqrt(h)/2 h^.5=1/(2 sqrt(h)) h + sqrt(h)/2 0=1/(2 sqrt(h)) x + sqrt(h)/2 x= -h

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

do you get it?